OpenStudy (anonymous):

Consider the problem stated below. “The perimeter of a rectangle is more than 36 cm. If the length of the rectangle is 3 more than the width, what are the minimum dimensions of the rectangle?” What is/are the unknown quantity(ies) in the problem?

OpenStudy (anonymous):

width? only?

OpenStudy (anonymous):

The unknown quantities are length, width, and perimeter. However, they're all related, as shown in the following equations: \[L = W+3\] \[P = 2(L+W)\] We can substitute using these equations to get an inequality: \[ P > 36 \] \[2((W+3)+W) > 36 \] \[4W + 6 > 36 \] \[W > \frac{15}{2} \] \[L > \frac{21}{2} \] Those are your minimum dimensions.

OpenStudy (anonymous):

Aooh! Thank youuuuuu.